OpenStudy (anonymous):

4(x+6)/3=2x-1 solve for x

OpenStudy (anonymous):

\[\frac{4(x+6)}{3}=2x-1\] \[4x+24=3(2x-1)\] \[4x+24=6x-3\] \[6x-4x=24+3\] \[2x=27\] \[x=\frac{27}{2}\]